# The wall confronting large language models

### v1

@@ 3 Scaling of Large Language Models
The convergence of modern ML schemes in general, and LLM’s in particular, is much harder to assess than for computer simulation methods, since they are based on a combination of deterministic and stochastic techniques in ultra-high-dimensional spaces with $d\gg 3$.
A milestone result in LLM’s research was the discovery that LLM’s not only survive overfitting (more parameters than data) but also seem to display “emergent” properties, i.e. they learn better as they get bigger. These scaling properties bear paramount technological implications, as they provide the basic booster for the relentless current big-tech race towards larger and larger LLMs, with parameters now extending well into the range of trillions. On a loose but telling note, this is still three decades short of the number of neural connections in the human brain, $10^{15}$, and yet they consume some one hundred million times more power (GWatts as compared to the very modest 20 Watts required by our brains).
@@ 3.1 Transformers as complex dynamical systems
This is not surprising, given that LLMs employ stochastic techniques with presumably strong non-Gaussian fluctuations, hence exposed to a much stronger Resilience of Uncertainty (RoU) than in the Gaussian world (for a precise definition of RoU, see Appendix I). The fact that LLM’s use stochastic techniques to predict the next word out of a pre-existing sequence is well known, but that does not prove that such stochastic processes are non-Gaussian. Very recent work shows that random rectified linear unit (ReLU) networks are indeed non-Gaussian processes [26]). One can, however, adduce both specific and general arguments to establish that this must indeed be the case. The specific one is that LLM architectures are centered upon transformers, whose basic action is to turn an input signal $x$ into an output $y$. The distinctive feature of transformers is that the number of neurons is the same across layers, which strongly facilitates the analogy with discrete dynamical systems. To highlight this analogy, it proves convenient to cast the transformer update as follows:
@@
A word of caution is in order. In the aforementioned discussion, we have assumed that the Loss function is equivalent to an approximation error. This association is not entirely correct because an approximation error equal to zero means perfect convergence to the desired target, whereas LLM practice informs us that pushing the loss function below a given threshold may be detrimental to the quality of the results (due to training).
@@
In this sense, the loss function is not a metric of the quality of an LLM’s predictions in the same way that an approximation error is in a “conventional” numerical scheme; instead, it is a pseudo-metric, which evades the convergence criteria for errors commonly adopted in scientific computing. There is nothing wrong with the loss per se, but given that it is not obvious that zero loss is the most desirable target, a decreasing trend in the loss function with increasing computational resources cannot necessarily be taken as an indication of success in its predictions. LLM loss is just a self-supervised learning signal. Slight mispredictions of the next word are often inconsequential, especially if they occur in intermediate reasoning traces that can be revised. In fact, driving the loss too low can lead to overfitting or mode collapse, and in rare cases a rising loss can be observed to correlate with improved model performance.
@@
As such, loss is a rather slippery and ill-defined concept whose lack of equivalence with convergence of errors would melt away were LLMs forced to perform quantitative scientific and mathematical tasks rather than being evaluated based on their qualitative alignment with meaning in a language context. A very serious scientific drawback of AI and LLMs in particular is the pace of their development, with the claims of awesome capability being superseded by new paradigms which are asserted to be far better before anyone has been able to make sense of how the original algorithm behaves. At the time of writing, large reasoning models (LRMs) are in the ascendent, said to be better still because they provide evidence of how they reach their answers — to say nothing of agentic AI which is supposed to perform better still by orchestrating the interactions between multiple LRMs. This article only considers LLMs. Indeed, through their design rules, LRMs actually render the loss function an even more tenuous measure of their performance. LLMs themselves are all of a kind: they are all based on the original transformer architecture dating back some eight years, and they all perform in similar ways, revealing a remarkably limited underlying diversity in design. Examples of their shared limitations include the fact that they are not very good at making jokes and have a extraordinarily limited repertoire of only two, while when asked to name a random number between 1 and 50, Anthropic’s, Google’s and OpenAI’s LLMs all come up with 27. xAI’s Grok produces 42 first, and then 27 [44]. In short, they lack novelty.
@@
For instance, it is not implausible to speculate that the ill-defined nature of the loss function as a metric of success plays a non-negligible role in the ubiquity of so-called “Potemkins” [^1] For the record, the term “Potemkins” stems from the Russian Grigory Potemkin, a field marshal and former lover of Empress Catherine II, who constructed phoney portable settlements solely to impress the Empress during her journey to Crimea in 1787. The structures would be disassembled after she passed, and re-assembled farther along her route where she could see them again., namely ”the illusion of understanding driven by answers irreconcilable with how any human would interpret a concept”, quoting from [23]. The authors of [23] go on to state: “We also find that these failures reflect not just incorrect understanding, but deeper internal incoherence in concept representations”. One can associate Potemkins with a general sort of “holographic accuracy”, meaning by this a convergence which appears to be satisfactory only as long as one does not start scratching beneath the surface. Holographic accuracy is possibly a general trait of modern society over and beyond LLM’s.
@@
To this end, the first observation is that in complex systems, and LLMs certainly qualify as such, non-Gaussian fluctuations are the rule rather than the exception. This is hardly a coincidence: non-Gaussian fluctuations feed the long-range spacetime correlations which lie at the very heart of complexity (telling it apart from pure chaos and randomness). This is a well-known hallmark of extended nonlinear systems far from equilibrium and plays a vital role in most dynamic structure formation processes [4]. It is therefore entirely reasonable to propose that such long-range correlations play a vital role in shaping the very structure of LLM’s (which, we recall are discrete dynamical systems), including their alleged “emergent properties” [33, 34, 35, 9].
@@ 7 Conclusions: an avoidable path to Degenerative AI
The considerations collected in this work configure a potential pathway to a doom-laden scenario which has been named Degenerative AI [17], meaning the catastrophic accumulation of self-feeding errors and inaccuracies, which appear particularly likely in the case of LLM’s trained on synthetic data.

### v2

@@ 3 Scaling of Large Language Models
The convergence of modern ML schemes in general, and LLMs in particular, is much harder to assess than for computer simulation methods, since they are based on a combination of deterministic and stochastic techniques in ultra-high-dimensional spaces with $d\gg 3$.
A milestone result in LLM research was the discovery that LLMs not only survive overfitting (more parameters than data) but also seem to display “emergent” properties, i.e. they learn better as they get bigger. These scaling properties bear paramount technological implications, as they provide the basic booster for the relentless current big-tech race towards larger and larger LLMs, with parameters now extending well into the range of trillions. On a loose but telling note, this is still three decades short of the number of neural connections in the human brain, $10^{15}$, and yet they consume some one hundred million times more power (GWatts as compared to the very modest 20 Watts required by our brains).
@@ 3.1 Transformers as complex dynamical systems
This is not surprising, given that LLMs employ stochastic techniques with presumably strong non-Gaussian fluctuations, hence exposed to a much stronger Resilience of Uncertainty (RoU) than in the Gaussian world (for a precise definition of RoU, see Appendix I). The fact that LLMs use stochastic techniques to predict the next word out of a pre-existing sequence is well known, but that does not prove that such stochastic processes are non-Gaussian. Very recent work shows that random rectified linear unit (ReLU) networks are indeed non-Gaussian processes [26]). One can, however, adduce both specific and general arguments to establish that this must indeed be the case. The specific one is that LLM architectures are centered upon transformers, whose basic action is to turn an input signal $x$ into an output $y$. The distinctive feature of transformers is that the number of neurons is the same across layers, which strongly facilitates the analogy with discrete dynamical systems. To highlight this analogy, it proves convenient to cast the transformer update as follows:
@@
A word of caution is in order. In the aforementioned discussion, we have assumed that the Loss function is equivalent to an approximation error. This association is not entirely correct because an approximation error equal to zero means perfect convergence to the desired target, whereas LLM practice informs us that pushing the loss function below a given threshold may be detrimental to the quality of the results (due to overtraining).
@@
In this sense, the loss function is not a metric of the quality of an LLM’s predictions in the same way that an approximation error is in a “conventional” numerical scheme; instead, it is a pseudo-metric, which evades the convergence criteria for errors commonly adopted in scientific computing. There is nothing wrong with the loss per se, but given that it is not obvious that zero loss is the most desirable target, a decreasing trend in the loss function with increasing computational resources cannot necessarily be taken as an indication of success in its predictions. LLM loss is just a self-supervised learning signal. Slight mispredictions of the next word are often inconsequential, especially if they occur in intermediate reasoning traces that can be revised. In fact, driving the loss too low can lead to overfitting or mode collapse.
@@
As such, loss is a rather slippery and ill-defined concept whose lack of equivalence with convergence of errors would melt away were LLMs forced to perform quantitative scientific and mathematical tasks rather than being evaluated based on their qualitative alignment with meaning in a language context. A very serious scientific drawback of AI and LLMs in particular is the pace of their development, with the claims of awesome capability being superseded by new paradigms which are asserted to be far better before anyone has been able to make sense of how the original algorithm behaves. At the time of writing, large reasoning models (LRMs) are in the ascendent, said to be better still because they provide evidence of how they reach their answers — to say nothing of agentic AI which is supposed to perform better still by orchestrating the interactions between multiple LRMs. This article only considers LLMs. Indeed, through their design rules, LRMs actually render the loss function an even more tenuous measure of their performance. LLMs themselves are all of a kind: they are all based on the original transformer architecture dating back some eight years, and they all perform in similar ways, revealing a remarkably limited underlying diversity in design. Examples of their shared limitations include the fact that they are not very good at making jokes and have an extraordinarily limited repertoire of only two, while when asked to name a random number between 1 and 50, Anthropic’s, Google’s and OpenAI’s LLMs all come up with 27. xAI’s Grok produces 42 first, and then 27 [44]. In short, they lack novelty.
@@
For instance, it is not implausible to speculate that the ill-defined nature of the loss function as a metric of success plays a non-negligible role in the ubiquity of so-called “Potemkins” [^1] For the record, the term “Potemkins” stems from the Russian Grigory Potemkin, a field marshal and former lover of Empress Catherine II, who constructed phoney portable settlements solely to impress the Empress during her journey to Crimea in 1787. The structures would be disassembled after she passed, and re-assembled farther along her route where she could see them again., namely ”the illusion of understanding driven by answers irreconcilable with how any human would interpret a concept”, quoting from [23]. The authors of [23] go on to state: “We also find that these failures reflect not just incorrect understanding, but deeper internal incoherence in concept representations”. One can associate Potemkins with a general sort of “holographic accuracy”, meaning by this a convergence which appears to be satisfactory only as long as one does not start scratching beneath the surface. Holographic accuracy is possibly a general trait of modern society over and beyond LLMs.
@@
To this end, the first observation is that in complex systems, and LLMs certainly qualify as such, non-Gaussian fluctuations are the rule rather than the exception. This is hardly a coincidence: non-Gaussian fluctuations feed the long-range spacetime correlations which lie at the very heart of complexity (telling it apart from pure chaos and randomness). This is a well-known hallmark of extended nonlinear systems far from equilibrium and plays a vital role in most dynamic structure formation processes [4]. It is therefore entirely reasonable to propose that such long-range correlations play a vital role in shaping the very structure of LLMs (which, we recall are discrete dynamical systems), including their alleged “emergent properties” [33, 34, 35, 9].
@@ 7 Conclusions: an avoidable path to Degenerative AI
The considerations collected in this work configure a potential pathway to a doom-laden scenario which has been named Degenerative AI [17], meaning the catastrophic accumulation of self-feeding errors and inaccuracies, which appear particularly likely in the case of LLMs trained on synthetic data.